# QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow

### v1

@@ 5.4 Remarks on the lattice Liouville theory from quantum Teichmuller theory
In the literature, it is known that there is a lattice integrable model that also recovers the Liouville CFT in the thermodynamic limit[50, *Faddeev:2000if, *Faddeev:2002ms, *2008arXiv0810.4352K].
The transfer matrix of the lattice model is constructed using operators in the quantum Teichmuller theory of Kashaev[178, *Kashaev:2000ku, 64]. In appendix E, we review the construction of these operators and explain the connection to the $U_{q}(SL(2,\mathbb{R}))$ $6j$-symbols.
@@ 5.5 Exact bulk dual at finite $N$ and universal semi-classical gravity dual at large $N$
In this section, we focus on the continuous spectrum (5.16), which describes theories related to pure gravity, for example the Liouville theory[1, 35]. What is the connection with general holographic CFTs? The answer to this question was first proposed in [63], and recently also emphasized in the discussion for the role played by ensemble averages in holography in terms of coarse-graining and the eigenstate thermalization hypothesis (ETH)[181, *PhysRevA.43.2046, 183, 184, *Stanford:2020wkf, *Pollack:2020gfa, 187, 188, 189].
@@
General irrational large $N$ holographic CFTs are expected to have a discrete spectrum instead of a continuous one that we studied in this section, and similarly for the associated topological defect lines. However, the general large $N$ holographic CFTs have a large gap and a dense spectrum for heavy states above the black hole threshold that is approximately continuous; see Fig. 21. The idea is that for individual 2D holographic CFTs, the heavy state dynamics of these theories can be universally approximated by the results outlined in this section. The coarse-grained result has a continuous spectrum (5.16) with Cardy density of states[162], and the heavy-state matrix elements are given by random matrices, from a generalized version of the ETH [181, *PhysRevA.43.2046, 63, 189]. These coarse-grained matrix elements involving heavy states can be obtained by studying the bootstrap equations involving identity in certain dual channels[190, 191, *Kusuki:2022wns, 193], and then applying an argument similar to [194] to the OPE coefficients, to extend the regime of validity in the large $N$ limit. The final result is universally captured by the quantum group $U_{q}(SL(2,\mathbb{R}))$, where quantum Teichmuller theory and geometries emerge. More details on the average over open structure coefficients in a CFT and its connection to the emergence of gravity will be presented in a forthcoming publication [195].
@@
In other words, even with the presence of additional fields and fine structures in a 2D CFT, the universality of heavy-state statistics at large $N$ ensures the universal emergence of a thermodynamic description with a 3D Einsteinian gravity sector in the bulk. States below the threshold will contribute to extra fields coupled to the gravity sector. For example, we can have worldline actions coupled to the Einstein-Hilbert action as in [196]. In this sense, gravity serves as a universal long-range entangled phase of matter, when there are infinitely many degrees of freedom. Building on this discussion, we emphasize the two distinct roles that Liouville theory plays. On one hand, it serves as an exact dual to pure gravity in a precise sense [25, 197, *Henneaux:2019sjx, 168, 1]. On the other hand, it functions as a universal approximation for generic 2D holographic theories when viewed through a coarse-grained lens [63, 190, 189, 187, 188].
@@
The above line of logic on coarse-graining explains why the bulk dual theory universally behaves as semi-classical gravity at large $N$. In terms of the RG operator and 3D quantum states, this is the statement that [^33] This is to be compared with an intrinsic ensemble of CFTs, which correspond to |$\ket{\Psi}=\sum_{i}p_{i}\ket{\Psi}_{i}$, where $p_{i}$ is a chosen weight for the ensemble average.:
@@ 6.1 Topological holographic principle
The topological holographic principle was proposed in [5, 7, *Kong:2020cie, *Kong:2020jne, *Chatterjee:2022kxb, 6, 11, 12, 13, 14, 15, 16, 17]. Practically, it states that the path integral of a $D$ dimensional theory with symmetry on a manifold $\Sigma_{D}$ can be expressed in terms of the path integral of a $D+1$ dimensional TQFT over a manifold $\Sigma_{D}\times I$, where $I$ is an interval. This manifold $\Sigma_{D}\times I$ is often referred to as a “sandwich”. The two boundaries of the sandwich are assigned different boundary conditions: one is the physical boundary condition, and the other is the topological boundary condition. In this paper, we associate the quantum state ${}_{\Lambda}\bra{\Omega}$ | with the physical boundary and the SymTFT ground state |$\ket{\Psi}_{\Lambda}$ with the topological boundary combined with the bulk.
@@ 6.2.1 Crossing symmetry and emergence of 3D bulk
Which of the ground states is selected in (4.10)? The information lies in the labels $\sigma$ in (4.10). In the CFT, $\sigma$ labels the conformal boundary conditions where in the collection at least one of them preserves one copy of the chiral algebra  [^34] As it is noted in [36], the framework does not require maximally extending the chiral algebra and allows one to include boundary conditions that break part of the chiral symmetries. The set of labels correspond to objects of a module category $\mathcal{M}_{\mathcal{C}}$ of ${\mathcal{C}}$. Each $\mathcal{M}_{\mathcal{C}}$ corresponds to a topological boundary condition of the Turaev-Viro TQFT[36, 199, 200, 14].
@@
The ground states of the Levin-Wen and Turaev-Viro models satisfy the no-flux constraint, and it has been shown in [27, 98] that this constraint ensures crossing symmetry and re-triangulation invariance, as expressed in (6.2). Conversely, wavefunctions derived from (6.2) inherently satisfy the no-flux condition, identifying them as ground states. Intuitively, this connection arises because the plaquette projection operator in the Levin-Wen model is constructed by effectively gluing a tetrahedron onto the wavefunctions. The work of FRS [36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw] demonstrates an equivalence between the associativity coefficients $\tilde{C}^{nmm^{\prime}}_{c_{1}c_{2}c_{3}}$ and the BCFT OPE coefficients, translating them into the crossing relation in the CFT. This identification is a direct manifestation of locality in the boundary-to-bulk map.
@@ 6.2.2 Modular invariance and anyon condensation/gauging
Now let’s explain the remaining conditions. As shown in FRS [36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw], to get a self-consistent CFT which is crossing symmetric and modular-invariant on arbitrary Riemann surfaces, we need to find a so-called symmetric special Frobenius algebra $\mathcal{A}\subset\mathcal{C}$, for $\mathcal{C}$ the Moore-Seiberg category which is a modular tensor category. A Frobenius algebra is a collection of objects in $\mathcal{C}$, together with a product structure that maps $\mathcal{A}\otimes\mathcal{A}\to\mathcal{A}$. The algebra admits modules/representations too, and they also form a fusion category $\mathcal{M}(\mathcal{A})$. Each representation $X\in\mathcal{M}(\mathcal{A})$ is also a collection of objects in $\mathcal{C}$. They are characterized by the action of the algebra on them. i.e. $\rho_{X}:\mathcal{A}\otimes X\to X$. $\rho_{X}$ is the counter-part of representation matrices of a group $G$ while $\mathcal{A}$ plays the role of the group $G$.
@@
The condensed collection of lines, which forms a Lagrangian algebra $\mathcal{L}_{\mathcal{A}}$ is known to be in 1-1 correspondence with modular invariants [212]. Therefore, each topological boundary $\mathcal{B}_{\mathcal{A}}$ controls the spectrum of local operators in the CFT in turn ensures that this spectrum is modular invariant. The condensation of lines can also be understood as the “gauging” of higher-form symmetries, and it has been proposed that the holographic duals of rational CFTs can be understood as taking the path integral of a pair of Chern-Simons theories and gauging the corresponding higher-form symmetries[153]. Before gauging, the ground states of the TQFT are on equal footing, and each produces some combination of CFT conformal blocks [94] that generically cannot exist independently in 2D (i.e. not modular invariant). After gauging, however, it singles out a unique ground state of the TQFT, so that the 3D bulk can be effectively collapsed, allowing the 2D theory to stand on its own, suggesting that the latter has an anomaly free spectrum. As an example, consider the theory defined on the torus. This gauging produces a unique choice of spectrum for the sum over the left and right moving sectors, ensuring modular invariance and intrinsic consistency as a 2D CFT. This pairing of left and right movers can also be visualised by unfolding the doubled Chern-Simons theory at the boundary, and turning the latter into an interface or surface operator inside a single copy of Chern-Simons theory[213, 200, 199], as shown in Fig. 23. In fact, this unfolded picture lies at the heart of the FRS construction [36].
@@
For example, diagonal RCFT corresponds to condensation of diagonal pairs of lines $i\otimes\bar{i}$, $\forall i\in\mathcal{C}$. This corresponds to taking $\mathcal{M}_{\mathcal{C}}=\mathcal{C}$, or equivalently, $\mathcal{A}=1$. This module category/Frobenius algebra exists in all fusion categories $\mathcal{C}$. Liouville theory corresponds to this diagonal gauging for two copies of the Virasoro TQFT, as explained in [1, 25], and illustrated in (5.28). Other choices of condensation can provide theories with states of non-zero integer spin [^35] A nontrivial Frobenius algebra $\mathcal{A}$ will provide a theory other than Liouville theory, whose bulk SymTFT is two copies of the Virasoro TQFT, potentially with non-zero spinning states.. We note that for more general irrational CFTs one would have to look for more non-trivial $\mathcal{A}$, which is certainly a highly non-trivial problem for irrational $\mathcal{C}$ [^36] Strictly speaking, since there are now infinitely many objects, this is no longer a fusion category. However, the success in Liouville theory suggests the existence of a generalized mathematical framework that extends to the irrational case..
@@
The bottom line is that each such Frobenius algebra $\mathcal{A}$ determines a modular invariant full CFT associated with the Moore-Seiberg data $\mathcal{C}$, and at the same time, the $\mathcal{M}(A)$ for given $\mathcal{A}$ can be identified with $\mathcal{M}_{\mathcal{C}}$[36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw]. In other words, the Frobenius algebra $\mathcal{A}$ determines both the full modular invariant CFT spectrum, and also its boundary conditions and boundary changing operators.
@@
In fact, for each individual theory, there can be multiple ways to realize it as a sandwich. As noted in the Introduction, the location where dynamics is encoded depends on the extent to which the theory’s symmetries are explicitly realized in the SymTFT. When the SymTFT captures only a limited subset of the CFT symmetries, the physical boundary state $\langle\Omega|$ retains a larger portion of the dynamical data. Conversely, as more symmetries are made explicit within the SymTFT, the physical boundary state simplifies accordingly. If all the symmetries are made explicit in the SymTFT, the physical boundary becomes purely kinematical, and the bulk together with the topological boundary, determines the theory’s dynamics.
@@
We have demonstrated that the exact correspondence between a 2D boundary and a 3D bulk theory arises from the full set of CFT self-consistency relations, which can be succinctly summarized by crossing symmetry and modular invariance [52, 60, 65, *Moore:1989vd, 36]. These self-consistency relations of 2D CFTs can also be elegantly encapsulated into algebraic equations connected to the 3D bulk SymTFT.
@@ 6.4.1 Wormholes, null states, and factorization
Recently, there has been growing interest in spacetime wormholes due to the remarkable observation made in [219, 220] that in some models, the inclusion of certain wormhole saddles in the Euclidean gravitational path integral can reproduce the Page curve at leading order in $G_{N}$, a result that has been proposed as a key feature for ensuring unitary black hole evaporation, resolving the information paradox [221, *Page:1993wv, *Page:1993df]. Although these calculations involve large $N$ approximations and do not offer a dynamical mechanism explaining how information escapes from black holes, they represent the first successful reproduction of the Page curve through a gravity computation. This is particularly notable because the problem has long been considered sensitive to UV dynamics, and an answer was not anticipated within the framework of semi-classical approximations. On the other hand, the inclusion of new wormhole saddles not only led to the factorization puzzle [224, *Yin:2007at, *Arkani-Hamed:2007cpn], but also calls into question the precise interpretation of the holographic correspondence. Should a given bulk describes a unique CFT, or an ensemble average of them [227, 218]? The modern perspective is that, in general, semi-classical gravity is a coarse-grained description of the dual chaotic CFT[227, 183, 196], in the sense of the ETH[181, *PhysRevA.43.2046], as we explained above.
@@
In fact, in the spirit of [220, 228], we might formulate the information paradox as the following problem: we have a matrix (reduced density matrix for radiation) with a large dimension, which is approximately diagonal (thermal density matrix[221]) , and this seems to contradict the rank of this matrix, bounded by the entropy $e^{S}$[229, *Hawking:1975vcx]. One resolution to this problem lies in the off-diagonal elements of the matrix, which reduces the rank of the matrix. The observation made in [220, 219] is that in certain computations such as the higher moments of the reduced density matrix of radiation, we can use the Euclidean spacetime wormhole saddles to reduce the rank, at leading orders in $G_{N}$. In similar settings, it is also shown that we can also reduce the rank of the Gram matrix of certain overcomplete black hole microstates constructed from dust particles, using wormhole saddles, to the black hole entropy $e^{A/4}$, at leading order in $G_{N}$[231, *Balasubramanian:2022lnw, *Climent:2024trz, *Geng:2024jmm].
@@
Is there a more microscopic and non-perturbative explanation for this reduction in rank? An answer to this question was proposed in [228], which highlighted that the reduction of rank in a matrix corresponds to the existence of kernel or null states in the theory, inherently tied to constraints and gauge equivalence in quantum gravity. In the current context, this gauge equivalence is suggested to arise from an equivalence relation imposed on geometries with different topologies, a perspective emphasized in [235] [^41] See also earlier works pointing out that the number of connected components of space cannot serve as a non-perturbative diffeomorphism-invariant operator in quantum gravity [236, *Bao:2015nca, *Berenstein:2016pcx, *Berenstein:2017abm]..
@@
Each choice of topological boundary condition appears to correspond to a choice of an “$\alpha$ state” in the language of [228, 241, *GIDDINGS1988854], where the partition functions factorize, as suggested by [153]. In our formalism, this is encoded in our state |$\ket{\Psi}_{\Lambda}$. However, as explained above, our perspective differs slightly from [153]. Specifically, we do not collapse the entire bulk and place the topological structure directly on top the physical boundaries, nor do we view the bulk as entirely trivial (which is the region beyond the topological boundary where anyon condensation occured). The bulk remains non-trivial within the sandwich, and in this light, the factorization puzzle can be resolved by identifying the quantum measure for the gravitational path integral over the non-trivial bulk which is encoded in the topological boundary condition, or the state |$\ket{\Psi}_{\Lambda}$.
@@
For example, translating Fig. 27 back into the semi-classical picture shown in Fig. 26, understanding the additional elements required for factorization in semi-classical gravity, and exploring connections to earlier proposals in other gravitational or string-theoretic models, such as those in [228, 243, *Blommaert:2021fob, 245, *Saad:2021uzi, 247, *Eberhardt:2020bgq]—are important directions for future investigation. Recently, it has also been proposed that an underlying symmetry principle, related to a gauge constraint, could eliminate wormhole contributions in the gravitational path integral, producing a unitary quantum theory. It is also worth understanding the precise relation between the Hopf algebra described in [249] [^43] Quantum groups are Hopf algebras by definition. In the case of rational CFTs, recovering a quantum group from the modular tensor category associated with the CFT can be done by the Tannaka duality., the “baby universe category” proposed in [250], and the Moore-Seiberg symmetry category $\mathcal{C}$ for 2D CFTs.

### v2

@@ 5.4 Remarks on the lattice Liouville theory from quantum Teichmuller theory
In the literature, it is known that there is a lattice integrable model that also recovers the Liouville CFT in the thermodynamic limit[42, *Faddeev:2000if, *Faddeev:2002ms, *2008arXiv0810.4352K].
The transfer matrix of the lattice model is constructed using operators in the quantum Teichmuller theory of Kashaev[174, *Kashaev:2000ku, 58]. In appendix E, we review the construction of these operators and explain the connection to the $U_{q}(SL(2,\mathbb{R}))$ $6j$-symbols.
@@ 5.5 Exact bulk dual at finite $N$ and universal semi-classical gravity dual at large $N$
In this section, we focus on the continuous spectrum (5.16), which describes theories related to pure gravity, for example the Liouville theory[1, 27]. What is the connection with general holographic CFTs? The answer to this question was first proposed in [57], and recently also emphasized in the discussion for the role played by ensemble averages in holography in terms of coarse-graining and the eigenstate thermalization hypothesis (ETH)[177, *PhysRevA.43.2046, 179, 180, *Stanford:2020wkf, *Pollack:2020gfa, 183, 184, 185].
@@
General irrational large $N$ holographic CFTs are expected to have a discrete spectrum instead of a continuous one that we studied in this section, and similarly for the associated topological defect lines. However, the general large $N$ holographic CFTs have a large gap and a dense spectrum for heavy states above the black hole threshold that is approximately continuous; see Fig. 21. The idea is that for individual 2D holographic CFTs, the heavy state dynamics of these theories can be universally approximated by the results outlined in this section. The coarse-grained result has a continuous spectrum (5.16) with Cardy density of states[158], and the heavy-state matrix elements are given by random matrices, from a generalized version of the ETH [177, *PhysRevA.43.2046, 57, 185]. These coarse-grained matrix elements involving heavy states can be obtained by studying the bootstrap equations involving identity in certain dual channels[186, 187, *Kusuki:2022wns, 189], and then applying an argument similar to [190] to the OPE coefficients, to extend the regime of validity in the large $N$ limit. The final result is universally captured by the quantum group $U_{q}(SL(2,\mathbb{R}))$, where quantum Teichmuller theory and geometries emerge. More details on the average over open structure coefficients in a CFT and its connection to the emergence of gravity will be presented in a forthcoming publication [191].
@@
In other words, even with the presence of additional fields and fine structures in a 2D CFT, the universality of heavy-state statistics at large $N$ ensures the universal emergence of a thermodynamic description with a 3D Einsteinian gravity sector in the bulk. States below the threshold will contribute to extra fields coupled to the gravity sector. For example, we can have worldline actions coupled to the Einstein-Hilbert action as in [192]. In this sense, gravity serves as a universal long-range entangled phase of matter, when there are infinitely many degrees of freedom.  [^34] In fact, since we propose that the bulk emerges from the Virasoro fusion kernels, including the light states related by analytic continuation, the entire CFT may be understood as arising from anyon condensation within a larger TQFT, where the Virasoro TQFT appears as a subsector.Building on this discussion, we emphasize the two distinct roles that Liouville theory plays. On one hand, it serves as an exact dual to pure gravity in a precise sense [17, 193, *Henneaux:2019sjx, 164, 1]. On the other hand, it functions as a universal approximation for generic 2D holographic theories when viewed through a coarse-grained lens [57, 186, 185, 183, 184].
@@
The above line of logic on coarse-graining explains why the bulk dual theory universally behaves as semi-classical gravity at large $N$. In terms of the RG operator and 3D quantum states, this is the statement that [^35] This should be compared with an intrinsic ensemble of CFTs (or equivalently Lagrangian algebras, as we will explain), which correspond to a state of the form |$\ket{\Psi}=\sum_{i}p_{i}\ket{\Psi}_{i}$, where $p_{i}$ represents the chosen weights for the ensemble average. This class of ket states is interpreted as having a bulk dual consisting only of smooth geometries, as studied in [195, 196, 197], and more recently in connection with SymTFT in [198, 199].:
@@ 6.1 Topological holographic principle
The topological holographic principle was proposed in [6, 7, *Kong:2020cie, *Kong:2020jne, *Chatterjee:2022kxb, 8, 9, 203, 204, 10, 54, 205, 94, 96]. Practically, it states that the path integral of a $D$ dimensional theory with symmetry on a manifold $\Sigma_{D}$ can be expressed in terms of the path integral of a $D+1$ dimensional TQFT over a manifold $\Sigma_{D}\times I$, where $I$ is an interval. This manifold $\Sigma_{D}\times I$ is often referred to as a “sandwich”. The two boundaries of the sandwich are assigned different boundary conditions: one is the physical boundary condition, and the other is the topological boundary condition. In this paper, we associate the quantum state ${}_{\Lambda}\bra{\Omega}$ | with the physical boundary and the SymTFT ground state |$\ket{\Psi}_{\Lambda}$ with the topological boundary combined with the bulk.
@@ 6.2.1 Crossing symmetry and emergence of 3D bulk
Which of the ground states is selected in (4.10)? The information lies in the labels $\sigma$ in (4.10). In the CFT, $\sigma$ labels the conformal boundary conditions where in the collection at least one of them preserves one copy of the chiral algebra.
The set of labels correspond to objects of a module category $\mathcal{M}_{\mathcal{C}}$ of ${\mathcal{C}}$. Each $\mathcal{M}_{\mathcal{C}}$ corresponds to a topological boundary condition of the Turaev-Viro TQFT[28, 206, 207, 54].
@@
The ground states of the Levin-Wen and Turaev-Viro models satisfy the no-flux constraint, and it has been shown in [19, 92] that this constraint ensures crossing symmetry and re-triangulation invariance, as expressed in (6.2). Conversely, wavefunctions derived from (6.2) inherently satisfy the no-flux condition, identifying them as ground states. Intuitively, this connection arises because the plaquette projection operator in the Levin-Wen model is constructed by effectively gluing a tetrahedron onto the wavefunctions. The work of FRS [28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw] demonstrates an equivalence between the associativity coefficients $\tilde{C}^{nmm^{\prime}}_{c_{1}c_{2}c_{3}}$ and the BCFT OPE coefficients, translating them into the crossing relation in the CFT. This identification is a direct manifestation of locality in the boundary-to-bulk map.
@@ 6.2.2 Modular invariance and anyon condensation/gauging
Now let’s explain the remaining conditions. As shown in FRS [28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw], to get a self-consistent CFT which is crossing symmetric and modular-invariant on arbitrary Riemann surfaces, we need to find a so-called symmetric special Frobenius algebra $\mathcal{A}\subset\mathcal{C}$, for $\mathcal{C}$ the Moore-Seiberg category which is a modular tensor category. A Frobenius algebra is a collection of objects in $\mathcal{C}$, together with a product structure that maps $\mathcal{A}\otimes\mathcal{A}\to\mathcal{A}$. The algebra admits modules/representations too, and they also form a fusion category $\mathcal{M}(\mathcal{A})$. Each representation $X\in\mathcal{M}(\mathcal{A})$ is also a collection of objects in $\mathcal{C}$. They are characterized by the action of the algebra on them. i.e. $\rho_{X}:\mathcal{A}\otimes X\to X$. $\rho_{X}$ is the counter-part of representation matrices of a group $G$ while $\mathcal{A}$ plays the role of the group $G$.
@@
The condensed collection of lines, which forms a Lagrangian algebra $\mathcal{L}_{\mathcal{A}}$ is known to be in 1-1 correspondence with modular invariants [219]. Therefore, each topological boundary $\mathcal{B}_{\mathcal{A}}$ controls the spectrum of local operators in the CFT in turn ensures that this spectrum is modular invariant. The condensation of lines can also be understood as the “gauging” of higher-form symmetries, and it has been proposed that the holographic duals of rational CFTs can be understood as taking the path integral of a pair of Chern-Simons theories and gauging the corresponding higher-form symmetries[149]. [^36] This idea dates back to [220, 221], and we thank Anatoly Dymarsky for bringing it to our attention. Before gauging, the ground states of the TQFT are on equal footing, and each produces some combination of CFT conformal blocks [72] that generically cannot exist independently in 2D (i.e. not modular invariant). After gauging, however, it singles out a unique ground state of the TQFT, so that the 3D bulk can be effectively collapsed, allowing the 2D theory to stand on its own, suggesting that the latter has an anomaly free spectrum. As an example, consider the theory defined on the torus. This gauging produces a unique choice of spectrum for the sum over the left and right moving sectors, ensuring modular invariance and intrinsic consistency as a 2D CFT. This pairing of left and right movers can also be visualised by unfolding the doubled Chern-Simons theory at the boundary, and turning the latter into an interface or surface operator inside a single copy of Chern-Simons theory[222, 207, 206], as shown in Fig. 23. In fact, this unfolded picture lies at the heart of the FRS construction [28].
@@
For example, diagonal RCFT corresponds to condensation of diagonal pairs of lines $i\otimes\bar{i}$, $\forall i\in\mathcal{C}$. This corresponds to taking $\mathcal{M}_{\mathcal{C}}=\mathcal{C}$, or equivalently, $\mathcal{A}=1$. This module category/Frobenius algebra exists in all fusion categories $\mathcal{C}$. Liouville theory corresponds to this diagonal gauging for two copies of the Virasoro TQFT, as explained in [1, 17], and illustrated in (5.28). Other choices of condensation can provide theories with states of non-zero integer spin [^37] A nontrivial Frobenius algebra $\mathcal{A}$ will provide a theory other than Liouville theory, whose bulk SymTFT is two copies of the Virasoro TQFT, potentially with non-zero spinning states.. We note that for more general irrational CFTs one would have to look for more non-trivial $\mathcal{A}$, which is certainly a highly non-trivial problem for irrational $\mathcal{C}$ [^38] Strictly speaking, since there are now infinitely many objects, this is no longer a fusion category. However, the success in Liouville theory and the existence of the Virasoro TQFT[192] suggests the existence of a generalized mathematical framework that extends to the irrational case..
@@
The bottom line is that each such Frobenius algebra $\mathcal{A}$ determines a modular invariant full CFT associated with the Moore-Seiberg data $\mathcal{C}$, and at the same time, the $\mathcal{M}(A)$ for given $\mathcal{A}$ can be identified with $\mathcal{M}_{\mathcal{C}}$[28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw]. In other words, the Frobenius algebra $\mathcal{A}$ determines both the full modular invariant CFT spectrum, and also its boundary conditions and boundary changing operators.
@@
In fact, for each individual theory, there can be multiple ways to realize it as a sandwich. As noted in the Introduction, the location where dynamics is encoded depends on the extent to which the theory’s symmetries are explicitly realized in the SymTFT. When the SymTFT captures only a limited subset of the CFT symmetries, the physical boundary state $\langle\Omega|$ retains a larger portion of the dynamical data. Conversely, as more symmetries are made explicit within the SymTFT, the physical boundary state simplifies accordingly. If all the symmetries are made explicit in the SymTFT, the physical boundary becomes purely kinematical, and the bulk together with the topological boundary, determines the theory’s dynamics. In 2D CFT, the corresponding maximal bulk theory arises from imposing all lines commuting with the Virasoro symmetry. If the theory possesses an extended chiral algebra, the bulk theory constructed from it will generally explicitly impose fewer symmetries, as the requirement that the lines commute with the full chiral algebra imposes additional constraints dictated by the extended algebra. [^41] We thank Conghuan Luo, Shu-Heng Shao and John McGreevy for the comments and discussions on this point. The collection of such topological lines is generally infinite, even in rational theories with an extended chiral algebra. In particular, when focusing only on the Virasoro algebra, the structure of many rational CFTs resembles that of irrational CFTs. In this light, we believe that many more theories may admit a non-trivial holographic bulk.
@@
We have demonstrated that the exact correspondence between a 2D boundary and a 3D bulk theory arises from the full set of CFT self-consistency relations, which can be succinctly summarized by crossing symmetry and modular invariance [44, 52, 59, *Moore:1989vd, 28]. These self-consistency relations of 2D CFTs can also be elegantly encapsulated into algebraic equations connected to the 3D bulk SymTFT.
@@ 6.4.1 Wormholes, null states, and factorization
Recently, there has been growing interest in spacetime wormholes due to the remarkable observation made in [228, 229] that in some models, the inclusion of certain wormhole saddles in the Euclidean gravitational path integral can reproduce the Page curve at leading order in $G_{N}$, a result that has been proposed as a key feature for ensuring unitary black hole evaporation, resolving the information paradox [230, *Page:1993wv, *Page:1993df]. Although these calculations involve large $N$ approximations and do not offer a dynamical mechanism explaining how information escapes from black holes, they represent the first successful reproduction of the Page curve through a gravity computation. This is particularly notable because the problem has long been considered sensitive to UV dynamics, and an answer was not anticipated within the framework of semi-classical approximations. On the other hand, the inclusion of new wormhole saddles not only led to the factorization puzzle [233, *Yin:2007at, *Arkani-Hamed:2007cpn], but also calls into question the precise interpretation of the holographic correspondence. Should a given bulk describes a unique CFT, or an ensemble average of them [236, 227]? The modern perspective is that, in general, semi-classical gravity is a coarse-grained description of the dual chaotic CFT[236, 179, 192], in the sense of the ETH[177, *PhysRevA.43.2046], as we explained above.
@@
In fact, in the spirit of [229, 237], we might formulate the information paradox as the following problem: we have a matrix (reduced density matrix for radiation) with a large dimension, which is approximately diagonal (thermal density matrix[230]) , and this seems to contradict the rank of this matrix, bounded by the entropy $e^{S}$[238, *Hawking:1975vcx]. One resolution to this problem lies in the off-diagonal elements of the matrix, which reduces the rank of the matrix. The observation made in [229, 228] is that in certain computations such as the higher moments of the reduced density matrix of radiation, we can use the Euclidean spacetime wormhole saddles to reduce the rank, at leading orders in $G_{N}$. In similar settings, it is also shown that we can also reduce the rank of the Gram matrix of certain overcomplete black hole microstates constructed from dust particles, using wormhole saddles, to the black hole entropy $e^{A/4}$, at leading order in $G_{N}$[240, *Balasubramanian:2022lnw, *Climent:2024trz, *Geng:2024jmm].
@@
Is there a more microscopic and non-perturbative explanation for this reduction in rank? An answer to this question was proposed in [237], which highlighted that the reduction of rank in a matrix corresponds to the existence of kernel or null states in the theory, inherently tied to constraints and gauge equivalence in quantum gravity. In the current context, this gauge equivalence is suggested to arise from an equivalence relation imposed on geometries with different topologies, a perspective emphasized in [244] [^44] See also earlier works pointing out that the number of connected components of space cannot serve as a non-perturbative diffeomorphism-invariant operator in quantum gravity [245, *Bao:2015nca, *Berenstein:2016pcx, *Berenstein:2017abm]..
@@
Each choice of topological boundary condition appears to correspond to a choice of an “$\alpha$ state” in the language of [237, 250, *GIDDINGS1988854], where the partition functions factorize, as suggested by [149]. In our formalism, this is encoded in our state |$\ket{\Psi}_{\Lambda}$. However, as explained above, our perspective differs slightly from [149]. Specifically, we do not collapse the entire bulk and place the topological structure directly on top the physical boundaries, nor do we view the bulk as entirely trivial (which is the region beyond the topological boundary where anyon condensation occured). The bulk remains non-trivial within the sandwich, and in this light, the factorization puzzle can be resolved by identifying the quantum measure for the gravitational path integral over the non-trivial bulk which is encoded in the topological boundary condition, or the state |$\ket{\Psi}_{\Lambda}$.
@@
For example, translating Fig. 27 back into the semi-classical picture shown in Fig. 26, understanding the additional elements required for factorization in semi-classical gravity, and exploring connections to earlier proposals in other gravitational or string-theoretic models, such as those in [237, 252, *Blommaert:2021fob, 254, *Saad:2021uzi, 256, *Eberhardt:2020bgq]—are important directions for future investigation. Recently, it has also been proposed that an underlying symmetry principle, related to a gauge constraint, could eliminate wormhole contributions in the gravitational path integral, producing a unitary quantum theory. It is also worth understanding the precise relation between the Hopf algebra described in [258] [^46] Quantum groups are Hopf algebras by definition. In the case of rational CFTs, recovering a quantum group from the modular tensor category associated with the CFT can be done by the Tannaka duality., the “baby universe category” proposed in [259], and the Moore-Seiberg symmetry category $\mathcal{C}$ for 2D CFTs.